OpenStudy (anonymous):

5-3-8-9-12-27-17-81-23-243-?

OpenStudy (dumbcow):

i dont get it?

OpenStudy (anonymous):

This one is simple There r two series going on 5+3=8+4=12+5=17+6=23 the other is 3x3=9x3=27x3=81 Is it clear???

OpenStudy (dumbcow):

ahh i always read too much into these. thanks anonda i hope that helped you so next number is 30